OpenStudy (anonymous):

where do i go from, 15(5x^2-1)-(x-2)

OpenStudy (anonymous):

What are you trying to do?

OpenStudy (anonymous):

factor the polynomial.

OpenStudy (anonymous):

egh. I'd start by getting it into the standard polynomial form

OpenStudy (anonymous):

the original is, 75x^3-150x^2-15x+30

OpenStudy (anonymous):

Ok, so factor a 15 out of everything.

OpenStudy (anonymous):

well i do factor by grouping

OpenStudy (anonymous):

I'd start by getting smaller numbers.

OpenStudy (anonymous):

well thats why i broke it up into 75x^3-150x^2 abd -15x+30

OpenStudy (anonymous):

and*

OpenStudy (anonymous):

\[75x^3-150x^2-15x+30\] \[= 15(5x^3 - 10x^2 - x + 2)\]

OpenStudy (anonymous):

Then regrouping and pulling a 5x from the first two terms you should have something nice.

OpenStudy (anonymous):

that way you would pull out a 75 and get 75x^2(x-2)15(x-2)

OpenStudy (anonymous):

so then its (75x^2-15) and (x-2) thenn 15(5x^2+1)(x-2)

OpenStudy (anonymous):

\[15(5x^3 - 10x^2 - x + 2) \] \[= 15[5x(x-2) -(x-2)]\] \[= 15(5x-1)(x-2)\]

OpenStudy (anonymous):

whoops i meant -1.

OpenStudy (anonymous):

so i basically had the right answer right?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Mine should be \[15[5x^2(x−2)−(x−2)]\] \[=15(5x^2-1)(x-2)\]

OpenStudy (anonymous):

gotcha

OpenStudy (anonymous):

alright i will be back in a few with some more